# When Fusion Helps and When It Breaks: View-Aligned Robustness in Same-Source Financial Imaging

### v1

@@ Abstract
We study same-source multi-view learning and adversarial robustness for next-day direction prediction with financial image representations. On Shanghai Gold Exchange (SGE) spot gold data (2005–2025), we construct two window-aligned views from each rolling window: an OHLCV-rendered price/volume chart and a technical-indicator matrix. To ensure reliable evaluation, we adopt leakage-resistant time-block splits with embargo and use Matthews correlation coefficient (MCC). We find that results depend strongly on the label-noise regime: we apply an ex-post minimum-movement filter that discards samples with realized $|r_{t+1}|<\tau$ to define evaluation subsets with reduced near-zero label ambiguity. This induces a non-monotonic data–noise trade-off that can reveal predictive signal but eventually increases variance as sample size shrinks; the filter is used for offline benchmark construction rather than an inference-time decision rule. In the stabilized subsets, fusion is regime dependent: early fusion by channel stacking can exhibit negative transfer, whereas late fusion with dual encoders and a fusion head provides the dominant clean-performance gains; cross-view consistency regularization has secondary, backbone-dependent effects. We further evaluate test-time $\ell_{\infty}$ perturbations using FGSM and PGD under two threat scenarios—view-constrained attacks that perturb one view and joint attacks that perturb both—and observe severe vulnerability at tiny budgets with strong view asymmetry. Late fusion consistently improves robustness under view-constrained attacks, but joint attacks remain challenging and can still cause substantial worst-case degradation.
@@ 1 Introduction
More views do not imply more robustness by default. Short-horizon financial prediction is notoriously difficult due to non-stationarity, heavy noise, and the ease with which evaluation can become overly optimistic under improper temporal splitting. At the same time, an increasingly common practical approach is to convert market time series into image-like representations—such as OHLCV/candlestick charts or technical-indicator matrices—so that convolutional models can exploit local spatial patterns and cross-variable structure. While such “financial vision” pipelines have reported promising results, two issues remain insufficiently clarified under a controlled and leakage-resistant evaluation: (i) when and why fusing multiple representations helps when the views are derived from the same underlying series, and (ii) how robust these single-view and fused predictors are under worst-case input perturbations, especially when the input consists of multiple aligned views.
This paper addresses these questions in a same-source multi-view setting for next-day direction prediction on Shanghai Gold Exchange (SGE) spot gold (2005–2025). From each rolling lookback window, we deterministically construct two aligned 2D views: an OHLCV-rendered price/volume view (ohlcv) and a technical-indicator matrix view (indic), both rendered to a common resolution and aligned to the same window end date. Unlike multi-source fusion where synchronization and provenance differences can confound results, our same-source design isolates the roles of representation choice and fusion strategy.
A key practical complication in next-day direction prediction is that labels can be dominated by near-zero moves, making the task effectively ambiguous and encouraging spurious correlations. To make this regime explicit, we introduce a minimum-movement threshold and construct evaluation subsets by discarding samples whose realized next-day return magnitude is below $\tau$. Importantly, this filtering uses the realized $r_{t+1}$ and is therefore an ex-post dataset construction step; it is used only to define label-noise regimes/subsets for controlled analysis and robustness benchmarking, not as a deployable rule available at prediction time. Within the two-view setting (both), we compare early fusion via channel stacking against late fusion via dual encoders with a fusion head, and further consider cross-view consistency regularization using temperature-scaled symmetric KL. All models are evaluated under a leakage-resistant time-block split with embargo to mitigate temporal leakage from overlapping windows (López de Prado, 2018), and we use Matthews correlation coefficient (MCC) as the primary metric.
Beyond clean accuracy, we examine adversarial robustness of these financial-image predictors. Small $\ell_{\infty}$-bounded perturbations are known to induce systematic failures in vision models (Goodfellow et al., 2015; Madry et al., 2018), and multi-branch architectures introduce additional threat-model choices: an adversary may perturb a single input stream or perturb all streams jointly. Using an explicit view–channel mapping, we define a view-aligned robustness protocol that supports (i) view-constrained attacks that perturb exactly one view and (ii) joint attacks that perturb both views, enabling fair comparisons between single-view baselines and two-view models (Yang et al., 2021). We emphasize that pixel-space perturbations on rendered views serve as a representation-level stress test of worst-case sensitivity for the chosen rendering and fusion design, rather than a claim about realistic perturbations in the underlying market or data-generation process.
Our experiments yield several practical findings. First, learnability hinges on label stabilization: without minimum-movement filtering (small $\tau$), most configurations hover around chance-level MCC, while moderate filtering unlocks predictive signal; overly strict filtering reduces effective sample size and increases variance, producing a non-monotonic data–noise trade-off. Second, fusion is regime dependent: early fusion can exhibit negative transfer under noisier operating points, whereas late fusion provides a more reliable default once labels stabilize and accounts for the main clean-performance gains. Third, naturally trained models are highly vulnerable under tiny $\ell_{\infty}$ budgets, and robustness is strongly view dependent: perturbing the indicator view can be particularly destructive for certain backbones. Finally, late fusion consistently improves robustness under view-constrained attacks by preserving an unperturbed evidence path, but coordinated joint attacks remain challenging and can still induce substantial worst-case degradation; diagnostics further suggest that consistency regularization primarily enforces logit-level alignment, while predictive signal is largely expressed through the feature-level fusion pathway rather than strong standalone branch classifiers.
Contributions.
  • A unified, leakage-resistant evaluation setup for same-source two-view prediction. Inspired by prior rendering schemes (e.g., Gao et al. (2022)), we build aligned ohlcv/indic views from SGE gold spot data and combine embargoed time-block splitting with MCC-based evaluation. To explicitly study label-noise regimes, we additionally report results under an ex-post minimum-movement subset definition (min_move) based on realized $|r_{t+1}|$.
  • Regime-dependent analysis of fusion under label stabilization. We show multi-view gains are regime dependent: early fusion may underperform in noisier regimes, while late fusion is a stable default once labels stabilize and accounts for the main clean-performance gains.
  • Multi-view threat scenarios and a view-aligned robustness protocol. We formalize view–channel mapping and evaluate both view-constrained attacks and joint $\ell_{\infty}$ attacks (FGSM/PGD) that perturb both views, enabling fair, view-aligned robustness comparisons to single-view baselines (cf. Yang et al. (2021)).
  • Diagnostics for late fusion and consistency. We report branch-level diagnostics (agreement, temperature-scaled symmetric KL, branch-wise vs. fused MCC) to interpret robustness trends and distinguish logit-level alignment from fused predictive power.
Organization.
We first describe the data, multi-view construction, task, and evaluation metrics, and then present the threat model and view-aligned robustness protocol in Section 4. Clean baselines and the role of label stabilization and fusion paradigms are reported in Section 6.1, followed by adversarial robustness results in Section 6.2. We conclude with discussion and implications in Section 7.
Code availability.
@@
Code will be released upon acceptance/publication at: https://github.com/mrui166/same-source-financial-imaging-robustness.
@@ 2.1 Financial Image Representations and Multi-View Fusion for Market Prediction
A common line of work in market prediction converts financial time series into 2D visual representations so that convolutional models can exploit local spatial patterns and cross-variable structures. Typical examples include candlestick-/OHLCV-style charts and chart-pattern based encodings, as well as indicator-driven matrices/heatmaps constructed from technical signals. Beyond ad-hoc chart rendering, a more principled “time-series-to-image” family encodes temporal structure into images such as Gramian Angular Fields and Markov Transition Fields, enabling multi-channel image learning from a single sequence (Wang and Oates, 2015). In financial settings, such imaging ideas have been instantiated in different ways, including technical-indicator image grids for trading (Sezer and Ozbayoglu, 2020a), bar-/chart-window image learners (Sezer and Ozbayoglu, 2020b; Jiang et al., 2023), and candlestick-pattern imaging for CNN-based recognition (Chen and Tsai, 2020). Related work also considers incorporating cross-asset market structure via graph-enhanced predictors, e.g., a graph convolutional feature based CNN for stock trend prediction (Chen et al., 2021). Recent surveys often group multi-source fusion methods for financial market prediction by when fusion is applied—pre-model (early/data), mid-model, or post-model (late/decision)—and note that predictive performance can be sensitive to this choice (Mehrmolaei and Saniee Abadeh, 2025). Related strands also explore learning from different representations of the same underlying data, sometimes fusing them (e.g., feature-level fusion in hybrid CNN–RNN pipelines) (Kim and Kim, 2019) or incorporating pattern-based representations such as candlesticks and sequence similarity (Liang et al., 2022).
@@
Within this landscape, our two-view construction follows the multi-view intuition that complementary representations can provide partially redundant evidence. We therefore treat complementarity as an empirical question and explicitly analyze regime dependence under label-stabilization settings (via a minimum-movement threshold). In particular, we adopt an OHLCV-rendered view that preserves price-action/volume cues and an indicator-matrix view that emphasizes trend/momentum signals; our same-source two-view design is inspired by prior multi-source image-integration approaches in financial forecasting (Gao et al., 2022), while our focus is on the same-source case to remove cross-source alignment confounds. We focus on a same-source setting where both views are deterministically derived from an aligned rolling window of a single market time series (SGE gold spot), which isolates the role of representation and fusion from confounding cross-source alignment issues.
@@
A key practical takeaway from prior fusion work is that multi-view learning is not automatically beneficial: gains depend on whether the views are complementary and on careful representation alignment and fusion design (Baltrusaitis et al., 2019; Rahate et al., 2022). More broadly, financial forecasting pipelines are also prone to subtle time leakage (e.g., through rolling normalization, window overlap, or hyperparameter search) and selection effects from repeated backtests; protocol designs such as purged CV/embargo and diagnostics such as PBO/DSR have been advocated to diagnose and quantify these risks (López de Prado, 2018; Bailey et al., 2017; Bailey and López de Prado, 2014). Related concerns about leakage undermining benchmark validity have also been discussed in other areas (e.g., benchmark reinforcement against memorization in LLM QA) (Fang et al., 2025), reinforcing the general point that evaluation reliability depends on making leakage assumptions explicit. Motivated by this, our study emphasizes a controlled comparison between early fusion (channel stacking) and late fusion (dual encoders with a fusion head), and interprets gains in an evaluation-reliability–oriented context of leakage-resistant time splits (Section 5.4) and label-stabilization regimes (Section 5.3). Many closely related financial-imaging papers report improvements under relatively controlled evaluation, whereas we center the analysis on evaluation protocol choices (e.g., leakage control, operating-point selection) and fusion paradigms as first-class factors that shape reliability.
@@ 2.2 Adversarial Robustness and Threat Models for Multi-View Inputs
Adversarial robustness has been extensively studied in vision, where small $\ell_{\infty}$-bounded perturbations can induce systematic misclassification (Szegedy et al., 2014; Goodfellow et al., 2015). Canonical first-order attacks include FGSM (Goodfellow et al., 2015) and multi-step PGD (Madry et al., 2018), which have become standard stress tests for probing worst-case sensitivity under first-order adversaries. Beyond generic vision benchmarks, a smaller but growing line of work considers adversarial sensitivity in financial vision models built on chart-like representations. For example, Chen et al. study adversarial robustness of a deep convolutional candlestick learner built on Gramian Angular Fields (GAF) encodings and propose perturbation sampling/adversarial training to improve stability under structured perturbations of candlestick inputs (Chen et al., 2020). These studies are complementary to ours: they focus on candlestick-pattern classification and representation-specific attacks/defenses, whereas we evaluate next-day direction prediction under leakage-resistant time splits and explicitly compare view-constrained vs. joint threat models for same-source two-view fusion. Beyond single-input settings, robustness questions also arise for multi-branch and multimodal models, where the threat model can be defined per input stream (single-source/view-constrained attacks), and where fusion architecture can modulate robustness and failure modes (Yang et al., 2021). In our setting, we evaluate an inference-time, white-box, untargeted FGSM/PGD threat model instantiated as raw pixel-space $\ell_{\infty}$-bounded perturbations on rendered views, with optional standardized-space execution via budget conversion that preserves the underlying raw-space constraint. We emphasize that this pixel-space threat model is a representation-level stress test of worst-case sensitivity for the chosen rendering and fusion design, rather than a claim about a realistic financial adversary operating in the underlying market or data-generation process.
Time-series settings introduce additional subtleties: perturbations can be defined either in the raw temporal domain or on downstream representations (e.g., rendered images), and the choice determines what “small” means and which invariances are being tested. Prior work has instantiated adversarial attacks directly on time-series classifiers and shown that standard gradient-based methods can transfer to temporal signals (Fawaz et al., 2019). In financial pipelines, where inputs often undergo deterministic preprocessing (windowing, indicator computation, normalization), robustness evaluation therefore benefits from threat models that explicitly account for where perturbations are applied and how multi-view inputs are aligned.
@@ 3.1 Same-Source Data (SGE Gold Spot)
Source. Shanghai Gold Exchange (SGE), spot gold; data retrieved via Tushare (Tushare, 2025).
Time range. 2005-09-26–2025-08-28.
Time zone normalization. All timestamps are converted from Shanghai time (UTC+8) to UTC.
@@ 3.2 Multi-View Construction
Motivated by the multi-view formulation in MODII (Gao et al., 2022), we construct two complementary 2D image views from a single market data source: (i) an OHLCV-based price/volume view and (ii) a technical-indicator view. This multi-representation design enables CNNs to capture local temporal patterns and cross-variable structures. Both views are constructed from the same reference day and aligned to the same 15-day lookback window.
OHLCV images.
Motivated by the OHLCV view used in MODII (Gao et al., 2022), we construct an OHLCV image view (denoted as ohlcv) from the open, high, low, and close prices together with trading volume within a fixed lookback window of length $L_{\mathrm{ohlcv}}=15$ trading days. Each image is rendered on a black canvas at a fixed resolution of $256\times 256$ pixels. The top region (75% of the height) depicts OHLC bars (high–low vertical strokes with open/close ticks), and we further highlight the intraday price interval by shading the open–close range in gray. This design is motivated by the interval-valued perspective that an interval observation can carry richer information (range and level/trend) than a point observation (Han et al., 2012). The bottom region (25% of the height) encodes volume as a bar chart, where volumes are normalized by a rolling percentile-based scaler (divided by the trailing 95th percentile computed from a fixed-length reference window) and clipped to $[0,1]$.
Indicator images.
To complement the price view with trend and momentum cues beyond raw OHLCV, we construct an indicator-image view (indic) by converting recent technical-indicator trajectories into an image-like matrix. We adopt a MODII-style construction pipeline (Gao et al., 2022), which is conceptually related to earlier time-series-to-image conversions for financial trading (Sezer and Ozbayoglu, 2020a). Specifically, we compute a fixed set of $N=15$ standard indicators over a lookback window of $L_{\mathrm{indic}}=15$ trading days. Indicator choices and parameters are fixed a priori; we do not search over indicator sets or tune parameters using validation/test performance. For each sample, we stack the most recent $L_{\mathrm{indic}}$ values of each indicator into a $N\times L_{\mathrm{indic}}$ grayscale matrix (indicators on the vertical axis, time on the horizontal axis), and apply per-row min–max scaling to $[0,255]$ within the current window. We then reorder indicator rows using a correlation-based ranking computed on the training split and keep this order fixed for validation/test. Finally, the $15\times 15$ matrix is upsampled with nearest-neighbor interpolation to a fixed $256\times 256$ canvas (to match the image encoder input size) and rendered as a grayscale image.

### v2

@@ Abstract
We study same-source multi-view learning and adversarial robustness for next-day direction prediction using two deterministic, window-aligned image views derived from the same time series: an OHLCV-rendered chart (ohlcv) and a technical-indicator matrix (indic). To control label ambiguity from near-zero moves, we use an ex-post minimum-movement threshold min_move ($\tau$) based on realized $|r_{t+1}|$, defining an offline benchmark on the subset $|r_{t+1}|\geq\tau$. Under leakage-resistant time-block splits with embargo, we compare early fusion (channel stacking) and dual-encoder late fusion with optional cross-branch consistency. We then evaluate pixel-space $\ell_{\infty}$ evasion attacks (FGSM/PGD) under view-constrained and joint threat models. We find that fusion is regime dependent: early fusion can suffer negative transfer under noisier settings, whereas late fusion is a more reliable default once labels stabilize. Robustness degrades sharply under tiny budgets with stable view-dependent vulnerabilities; late fusion often helps under view-constrained attacks, but joint perturbations remain challenging.
keywords:
multi-view learning , same-source , fusion , adversarial robustness , financial imaging , view-aligned evaluation
[…]
[inst1]organization=Xingchong (Xi’an) Technology Co., Ltd., addressline=R&D Department, city=Xi’an, state=Shaanxi, country=China
@@ 1 Introduction
More views do not automatically translate into better robustness. Financial time series prediction is notoriously sensitive to label ambiguity, non-stationarity, and evaluation leakage. In particular, next-day direction labels can be dominated by near-zero moves, where the notion of “up” versus “down” is weakly defined and easily overwhelmed by noise. At the same time, there is growing interest in vision-style representations of market data—candlestick charts, indicator heatmaps, and other renderings—because they enable the reuse of mature computer-vision architectures and data-augmentation pipelines. However, when multiple views or modalities are involved, practical comparisons are often confounded by cross-source alignment issues, inconsistent timestamps, or missingness patterns, making it difficult to attribute gains to fusion design rather than data integration.
This work studies same-source multi-view learning and adversarial robustness in a controlled setting. We construct two deterministic, window-aligned image views from the same underlying time series: (i) an OHLCV-rendered price/volume chart (ohlcv) and (ii) a technical-indicator matrix view (indic). Because both views are derived from identical timestamps and a shared window definition, the setup isolates representation and fusion effects from cross-source alignment confounds. We benchmark standard convolutional vision backbones and compare early fusion (channel stacking) against dual-encoder late fusion with a fusion head, optionally augmented with a cross-branch consistency regularizer.
A key methodological issue is that the unconditional next-day direction task can be nearly unlearnable when labels are dominated by micro-moves. To study learnability and robustness under controlled label-ambiguity regimes, we use an ex-post minimum-movement threshold min_move ($\tau$) defined by the realized next-day return magnitude $|r_{t+1}|$. This label-dependent filtering changes the task from unconditional direction prediction to direction prediction conditional on $|r_{t+1}|\geq\tau$; it is used solely for offline benchmarking and is not an inference-time decision rule. To avoid leakage and keep comparisons controlled, we apply filtering once prior to splitting and then run the embargoed time-block split on the filtered index set; details are in Appendices˜D and D.
We further evaluate adversarial robustness under an explicit view–channel threat model. Using pixel-space $\ell_{\infty}$ attacks on the rendered inputs (FGSM/PGD) [1, 2], we distinguish between view-constrained attacks that perturb exactly one view and joint attacks that perturb both views simultaneously. This protocol makes it possible to ask whether multi-view learning provides robustness through redundancy when only one view is corrupted, and whether such redundancy persists under coordinated perturbations. We also conduct operating-point sensitivity checks (e.g., at $\tau=0.008$) to assess which qualitative findings are stable and which are operating-point dependent (Appendix˜I).
Our main findings, summarized in Section˜6.2 and discussed in Section˜7, can be distilled as follows: (i) fusion behavior is strongly regime dependent—early fusion can exhibit negative transfer under noisier settings, while late fusion is a more reliable default once labels stabilize; (ii) naturally trained models degrade sharply under tiny pixel-space budgets, and robustness is markedly view dependent, revealing stable view asymmetries; (iii) late fusion often degrades more gracefully under view-constrained attacks, but joint perturbations remain challenging; and (iv) consistency regularization primarily improves cross-branch alignment, while its effects on clean performance and robustness are conditional and can shift with $\tau$.
Contributions
We make the following contributions:
  • 1.
    Controlled same-source multi-view benchmark. We define a window-aligned dual-view imaging protocol for SGE gold spot data (2005–2025), enabling clean attribution of fusion effects without cross-source alignment confounds.
  • 2.
    Operating-point-aware evaluation. We use min_move to construct controlled label-ambiguity regimes for offline benchmarking, paired with leakage-resistant time-block splits with embargo to reduce temporal leakage (Sections˜5.4 and D).
  • 3.
    View–channel adversarial robustness analysis. We evaluate FGSM/PGD under view-constrained and joint threat models, characterize sharp degradation under tiny budgets, and document stable view-dependent vulnerabilities and the conditional benefits of late fusion (Section˜6.2).
Finally, we note the scope of this study: our baselines focus on convolutional vision backbones (Lite-CNN and ResNet18-P) and pixel-space attacks on rendered inputs. Extending the benchmark to Transformer-based architectures and broader threat models remains an important direction for future work.
Organization
After reviewing related work, we introduce the same-source multi-view setup, task, and evaluation protocol, and then specify the threat model and view-aligned robustness evaluation in Section˜4. Clean baselines and fusion behavior are reported in Section˜6.1, followed by adversarial robustness results in Section˜6.2. We discuss implications and limitations in Section˜7 and conclude in Section˜8. Additional implementation details and extended analyses—including the $2/255$ stress test, operating-point sensitivity at $\tau=0.008$, and full branch diagnostics—are provided in Appendices˜J, I and H.
Code availability
@@
Code to reproduce the experiments will be released upon acceptance at: https://github.com/mrui166/same-source-financial-imaging-robustness.
@@ 2.1 Financial Image Representations and Multi-View Fusion for Market Prediction
A common line of work in market prediction converts financial time series into 2D visual representations so that convolutional models can exploit local spatial patterns and cross-variable structures. Typical examples include candlestick-/OHLCV-style charts and chart-pattern based encodings, as well as indicator-driven matrices/heatmaps constructed from technical signals. Beyond ad-hoc chart rendering, a more principled “time-series-to-image” family encodes temporal structure into images such as Gramian Angular Fields and Markov Transition Fields, enabling multi-channel image learning from a single sequence [3]. In financial settings, such imaging ideas have been instantiated in different ways, including technical-indicator image grids for trading [4], bar-/chart-window image learners [4, 5], and candlestick-pattern imaging for CNN-based recognition [6]. Related work also considers incorporating cross-asset market structure via graph-enhanced predictors, e.g., a graph convolutional feature based CNN for stock trend prediction [7]. Recent surveys often group multi-source fusion methods for financial market prediction by when fusion is applied—pre-model (early/data), mid-model, or post-model (late/decision)—and note that predictive performance can be sensitive to this choice [8]. Related strands also explore learning from different representations of the same underlying data, sometimes fusing them (e.g., feature-level fusion in hybrid CNN–RNN pipelines) [9] or incorporating pattern-based representations such as candlesticks and sequence similarity [10].
@@
Within this landscape, our two-view construction follows the multi-view intuition that complementary representations can provide partially redundant evidence. We therefore treat complementarity as an empirical question and explicitly analyze regime dependence under label-stabilization settings (via a minimum-movement threshold). In particular, we adopt an OHLCV-rendered view that preserves price-action/volume cues and an indicator-matrix view that emphasizes trend/momentum signals; our same-source two-view design is inspired by prior multi-source image-integration approaches in financial forecasting [11], while our focus is on the same-source case to remove cross-source alignment confounds. We focus on a same-source setting where both views are deterministically derived from an aligned rolling window of a single market time series (SGE gold spot), which isolates the role of representation and fusion from confounding cross-source alignment issues.
@@
A key practical takeaway from prior fusion work is that multi-view learning is not automatically beneficial: gains depend on whether the views are complementary and on careful representation alignment and fusion design [12, 13]. More broadly, financial forecasting pipelines are also prone to subtle time leakage (e.g., through rolling normalization, window overlap, or hyperparameter search) and selection effects from repeated backtests; protocol designs such as purged CV/embargo and diagnostics such as PBO/DSR have been advocated to diagnose and quantify these risks [14, 15, 16]. Related concerns about leakage undermining benchmark validity have also been discussed in other areas (e.g., efforts to mitigate memorization-driven leakage in LLM QA benchmarks) [17], which serves as a reminder that evaluation reliability depends on stating leakage-related assumptions explicitly. Motivated by this, our study emphasizes a controlled comparison between early fusion (channel stacking) and late fusion (dual encoders with a fusion head), and interprets gains in an evaluation-reliability–oriented context of leakage-resistant time splits (Section˜5.4) and label-stabilization regimes (Section˜5.3). Many closely related financial-imaging papers report improvements under relatively controlled evaluation, whereas we center the analysis on evaluation protocol choices (e.g., leakage control, operating-point selection) and fusion paradigms as first-class factors that shape reliability.
@@ 2.2 Adversarial Robustness and Threat Models for Multi-View Inputs
Adversarial robustness has been extensively studied in vision, where small $\ell_{\infty}$-bounded perturbations can induce systematic misclassification [18, 1]. Canonical first-order attacks include FGSM [1] and multi-step PGD [2], which have become standard stress tests for probing worst-case sensitivity under first-order adversaries. Beyond generic vision benchmarks, a smaller but growing line of work considers adversarial sensitivity in financial vision models built on chart-like representations. For example, Chen et al. study adversarial robustness of a deep convolutional candlestick learner built on Gramian Angular Fields (GAF) encodings and propose perturbation sampling/adversarial training to improve stability under structured perturbations of candlestick inputs [19]. These studies are complementary to ours: they focus on candlestick-pattern classification and representation-specific attacks/defenses, whereas we evaluate next-day direction prediction under leakage-resistant time splits and explicitly compare view-constrained vs. joint threat models for same-source two-view fusion. Beyond single-input settings, robustness questions also arise for multi-branch and multimodal models, where the threat model can be defined per input stream (single-source/view-constrained attacks), and where fusion architecture can modulate robustness and failure modes [20]. In our setting, we evaluate an inference-time, white-box, untargeted FGSM/PGD threat model instantiated as raw pixel-space $\ell_{\infty}$-bounded perturbations on rendered views, with optional standardized-space execution via budget conversion that preserves the underlying raw-space constraint. We emphasize that this pixel-space threat model is a representation-level stress test of worst-case sensitivity for the chosen rendering and fusion design, rather than a claim about a realistic financial adversary operating in the underlying market or data-generation process.
Time-series settings introduce additional subtleties: perturbations can be defined either in the raw temporal domain or on downstream representations (e.g., rendered images), and the choice determines what “small” means and which invariances are being tested. Prior work has instantiated adversarial attacks directly on time-series classifiers and shown that standard gradient-based methods can transfer to temporal signals [21]. In financial pipelines, where inputs often undergo deterministic preprocessing (windowing, indicator computation, normalization), robustness evaluation therefore benefits from threat models that explicitly account for where perturbations are applied and how multi-view inputs are aligned.
@@ 3.1 Same-Source Data (SGE Gold Spot)
Source. Shanghai Gold Exchange (SGE), spot gold; data retrieved via Tushare [22].
Time range. 2005-09-26–2025-08-28 (all dates reported in UTC).
@@ 3.2 Multi-View Construction
Motivated by the multi-view formulation in MODII [11], we construct two complementary 2D image views from the same market-data window: (i) an OHLCV-based price/volume view (ohlcv) and (ii) a technical-indicator view (indic). Both views are generated from the same reference day and aligned to the same 15-day lookback window, enabling controlled comparisons across single-view and two-view settings. Our construction follows a MODII-style pipeline [11] and is conceptually related to prior time-series-to-image transformations for financial prediction [4]. Rendering, scaling, and ordering details are provided in B.
OHLCV view (ohlcv)
We render OHLC bars with a volume subpanel to capture recent level, range, and trading activity in a compact visual representation.
Indicator view (indic)
We compute a fixed set of $N=15$ standard indicators over the same lookback window and convert the resulting indicator-by-time matrix into a grayscale image; the indicator set and parameters are fixed a priori (B).